OpenStudy (anonymous):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y = Seven divided by x squared. + 10

OpenStudy (anonymous):

Is it \[\frac{ 7 }{ x^2 } + 10\] Or \[(\frac{ 7 }{ x })^2 +10\]

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Well I guess this could be one solution:\[g(x) = x^2\] \[f(x) = \frac{ 7 }{ x } + 10\]

OpenStudy (anonymous):

So \[f(g(x)) = \frac{ 7 }{ x^2 } + 10\]